OpenStudy (anonymous):

Find two consecutive odd integers the sum of whose square is 130. x+(x+2)=130 x+(x+2)-130=0 does this look right so far the square part kind of confuses me :(

hartnn (hartnn):

sum of squares means, x^2 + (x+2)^2 =130

hartnn (hartnn):

x^2 + (x+2)^2 =130 can u solve this ahead ?

OpenStudy (anonymous):

is it x^2(x+2)(x+2)-130=0 and then after you foil the quantity of (x+2)^2 then you just factor right?

hartnn (hartnn):

yes, but u missed a + sign in between, its x^2+(x+2)(x+2)-130=0 then foil.

hartnn (hartnn):

or u can use \((x+a)^2=x^2+2ax+a^2\)

OpenStudy (anonymous):

i would make a guess and see if it worked i guess 9 and 11 \[9^2+11^2=202\]nope too big

OpenStudy (anonymous):

guess i will try 7 and 9 yup, that works

OpenStudy (anonymous):

no need to solve some silly equation for this, a little experimentation will do the numbers are not that big, so you don't have too many choices

OpenStudy (anonymous):

when i foiled i got x^2+(x^2+2x+2x+4)-130=0 so then i would get x^2+x^2+4x+4-130=0 2x^4+4x-126=0 then i got lost after i did this...

hartnn (hartnn):

now divide the whole equation by 2, and u get a nice quadratic in x with leading co-efficient =1.

hartnn (hartnn):

and which you can factor easily.

OpenStudy (anonymous):

yeah i got thanks so much! :D

hartnn (hartnn):

ok, welcome ^_^